Based on data from a car bumper sticker​ study, when a car is randomly​ selected, the number of bumper stickers and the corresponding probabilities are as shown below.

x   P(x)
0   0.800
1   0.088
2   0.045
3   0.017
4   0.014
5   0.012
6   0.009
7   0.007
8   0.006
9   0.002

A. Does the given information describe a probability​ distribution?

B. Assuming that a probability distribution is​ described, find its mean and standard deviation.

C. Use the range rule of thumb to identify the range of values for usual numbers of bumper stickers.

D. Is it unusual for a car to have more than one bumper​ sticker? Explain.

a. No, because the probability of more than 1 bumper sticker is 0.119, which is greater than 0.5.

b. Yes, because the probabilities for random variables x from 2 to 9 are all less than 0.05.

c. No, because the probability of having 1 bumper sticker is 0.099, which is greater than 0.05.

d. Not enough information is given.
Answer #1

a)Yes as sum of all probability is equal to 1 and individual probability is between 0 and 1

b)

x f(x) xP(x) x2P(x)
0 0.800 0.000 0.000
1 0.088 0.088 0.088
2 0.045 0.090 0.180
3 0.017 0.051 0.153
4 0.014 0.056 0.224
5 0.012 0.060 0.300
6 0.009 0.054 0.324
7 0.007 0.049 0.343
7 0.006 0.042 0.294
9 0.002 0.018 0.162
total 0.508 2.068
E(x) =μ= ΣxP(x) = 0.5080
E(x2) = Σx2P(x) = 2.0680
Var(x)=σ2 = E(x2)-(E(x))2= 1.8099
std deviation=         σ= √σ2 = 1.3453

from above mean =0.508

and std deviation =1.3453

c)

minimum usual value =μ-2*σ = -2.183
maximum usual value =μ+2*σ = 3.199